OpenStudy (anonymous):

Essay; show all work. Consider that the age, x, of a unicorn in human equivalent years can be given by the formula f(x) = - 0.0021718x4 + 0.027326x3 – 0.3367x2 + 12.46x + 2.914. When a unicorn is 2.5 years old, what is its age in human equivalent years? What about when it is 12 years old?

OpenStudy (agreene):

You need to plug in the values at the end of the question for the x in the equation. Just plug it in a calculator and you're good to go.

OpenStudy (anonymous):

i worked it out already i just wanted to double check to make sure i have it correct

OpenStudy (anonymous):

f(x) = - 0.0021718x4 + 0.027326x3 – 0.3367x2 + 12.46x + 2.914 f(2.5) = -0.0021718(2.5)^4 + 0.027326(2.5)^3 – 0.3367(2.5)^2 + 12.46(2.5) + 2.914 = - 0.0848359375 + 0.42696875 – 2.104375 + 31.15 + 2.914 = 32,3017578125 f(12) = -0.0021718(12)^4 + 0.027326(12)^3 – 0.3367(12)^2 + 12.46(12) + 2.914 f(12) = -45.0344448 + 47.219328 - 48.4848 + 149.52 + 2.914 = 106.1340832

OpenStudy (anonymous):

Yes, so for example: f(2.5) = - 0.0021718*2.5^4 + 0.027326*2.5^3 – 0.3367*2.5^2 + 12.46*2.5 + 2.914

OpenStudy (agreene):

32.3017578125 and 106.1340832 are what I get.

OpenStudy (anonymous):

okay thanks guys am second guessing myself can i ask you to look at another problem for me i am stuck